For the following pairs of supply and demand equations, where x represents the quantity

demanded in units of a thousand and p is the unit price in dollars, find the equilbrium point.

(a) 0 = −2x

2 − p+80 and 0 = 15x− p+30

(b) 11p+3x = 66 and 2p

2 + p−x = 10


1
Expert's answer
2021-01-19T07:21:44-0500

a)


"-2x^2-p+80=0"

"15x-p+30=0"

"-2x^2-15x+50=0"

"2x^2+15x-50=0"

"x=2.5"

"p=67.5"

b)


"11p+3x=66"

"2p^2+p-x=10"


"3p^2+7p-48=0"

"x=11"

"p=3"